3) Is the reaction 2 NH, + C02-NILCONH, + H,O (dG--/3. 6 h/) spontaneous at 25°C if PS-0.08 atm, Pco 0.3 atm, and [NH,CONH, 1.8 M?
Answer #1

For the reaction to be spontaneous, Delta G of the reaction must be negative

K = \frac{[Products]}{[Reactants]} = \frac{[P_{NH_2CONH_2}]}{[P_{NH_3}]^2[P_{CO_2}]} = \frac{1.8}{(0.08)^2(0.3)} = 937.5

Using the formula

\Delta G = \Delta G_0 + RTln(K)

\Delta G = -13600 + (8.314)(273+25)ln(937.5) = -13600 + 16954.56 = 3354.562J = 3.354 kJ

Since Delta G is positive, hence the reaction is not spontaneous at this temperature and these pressures